OpenStudy (anonymous):

Given temperatures in degrees Fahrenheit, Celcius, or Kelvin, use the appropriate formula to convert to one of the other temperature units a. Celcius to Fahrenheit b. Fahrenheit to Celcius F = (9/5)C + 32 c. Celcius to Kelvin d. Kelvin to Celcius C = (5/9)(F – 32) e. Fahrenheit to Kelvin f. Kelvin to Fahrenheit K = C + 273.15

thomaster (thomaster):

What's up with a c and e? there are no formulas given

OpenStudy (anonymous):

they are the same as the ones under them

thomaster (thomaster):

So what's the question?

OpenStudy (anonymous):

how do you convert them and what do they come out to

OpenStudy (aravindg):

This might help: http://www.mathsisfun.com/temperature-conversion.html

thomaster (thomaster):

well fahrenheit to celcius is F = (9/5)C + 32 So to rewrite the formula to celcius to fahrenheit , you'd have to isolate C from the formula. First subtract 32 from both sides F-32=(9/5)C Then divide both sides by (9/5) So you'll get C = (f-32)/(9/5)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you know how to do the rest of them

thomaster (thomaster):

Well celcius to kelvin is K=C+273.15 simple as that Kelvin to celcius is just the other way around: C = K-273.16

thomaster (thomaster):

Fahrenheit to kelvin -> K = (F + 459.67) / (9/5) Kelvin to fahrenheit -> F = K * (9/5) - 459.67

OpenStudy (anonymous):

thank you do you mind helping me with some more questions

OpenStudy (anonymous):

Given quantities of energy in Calories, calories, or joules, use the appropriate formula to convert to one of the other energy units. a. calories to Calories b. Calories to calories 1 cal = 4.184 J c. Joules to calories d. calories to Joules 1 Cal = 1000 cal (note difference)

thomaster (thomaster):

calories is the same as Calories right? I think you mean kcal (kilo calories) kcal = 1000 cal 1 cal = 1/1000 kcal 1 cal = 4.184 Joules 1 Joule = 1/4.184 cal

OpenStudy (anonymous):

thank you can you do like 5 or 6 more 3. Given a Table of Specific Heats, compare the specific heats of substances to determine which substance takes the least or most energy to heat up. Delta(T)=T-T

thomaster (thomaster):

This is something completely different. You should post it in a new question.

OpenStudy (anonymous):

oh really sorry i don't understand this stuff